\begin{thebibliography}{99}
\bibitem{BPL2} N.Bebiano, R.Lemos and J.da Provid\^encia Jr., Inequalities for quantum relative entropy,   Linear Alg. Appl., Vol.401, pp.159-172 (2005).
\bibitem{BPL} N.Bebiano, J.da Provid\^encia Jr. and R.Lemos, Matrix inequalities in statistical mechanics, Linear Alg. Appl., Vol.376, pp.265-273 (2004).
%\bibitem{Bha} R.Bhatia, Matrix Analysis, Springer, 1997.
\bibitem{Bou} J.-C.Bourin, Some inequalities for norms on matrices and operators, Linear Alg.Appl., Vol.292, pp.139-154 (1999).
\bibitem{Fujii}  J.I.Fujii, A trace inequality arising from quantum information theory, Linear Alg. Appl., Vol.400, pp.141-146 (2005).
\bibitem{FYK}    S.Furuichi, K.Yanagi and K.Kuriyama, Fundamental properties of Tsallis relative entropy, 
                   J.Math.Phys.,Vol.45,pp.4868-4877 (2004).
\bibitem{FYK2} S.Furuichi, K.Yanagi and K.Kuriyama, A note on operator inequalities of Tsallis relative opeartor entropy,  Linear Alg. Appl.,Vol.407,pp.19-31 (2005).
\bibitem{Furuta} T.Furuta, Convergence of logarithmic trace inequalities via generalized Lie-Trotter formulae, 
                   Linear Alg. Appl., Vol.396, pp.353-372 (2005).
\bibitem{Gol} S.Golden, Lower bounds for the Helmholtz function, Phys. Rev., Vol.137, pp.B1127-B1128 (1965). 
\bibitem{Hei} E.Heinz, Beitr\"age zur St\"orungstheorie der Spektralzerlegung, Math. Ann., Vol.123, pp.415-438 (1951).
\bibitem{HP2}    F.Hiai and D.Petz, The Golden-Thompson trace inequality is complemented, Linear Alg. Appl.,Vol.181,pp.153-185 (1993).
\bibitem{HH} Y. Hong and R. A. Horn, The Jordan canonical form of a product of a Hermitian and a positive semidefinite matrix, Linear Alg. Appl.,Vol.147,pp.373-386(1991).
\bibitem{Kos}    H.Kosaki, Relative entropy for states: a variational expression, J.Operator Theory, Vol.16, pp.335-348 (1986).
\bibitem{LT} E.H. Lieb and W. Thirring, Inequalities for the moments of the eigenvalues of the Schrodinger Hamiltonian and their relation to Sobolev inequalities, 
in E.Lieb, B.Simon and A.S.Wightman eds., \lq\lq Studies in Mathematical Physics. Essays in Honor of Valentine Bargmann\rq\rq, Princeton University Press, Princeton, NJ, pp.269-303(1976).
\bibitem{Low} K.L\"owner, \"Uber monotone Matrixfunktionen, Math.Z., Vol.38, pp.177-216 (1934).

\bibitem{OP}     M.Ohya and D.Petz, Quantum Entropy and its Use, Springer-Verlag,1993.
%\bibitem{Araki}  H.Araki, On an inequality of Lieb and Thirring, Lett.Math.Phys.,Vol.19,pp.167-170 (1990).
\bibitem{Ped} G.K.Pedersen, Some operator monotone functions, Proc. Amer. Math. Soc., Vol.36, pp.309-310 (1972).
\bibitem{Petz1}  D.Petz, A survey of certain trace inequalities, in Functional Analysis and Operator Theory, pp.287-298, 
                   Banach Center Publications, 30 (Warszawa 1994).
\bibitem{Petz2}  D.Petz, A variational expression for the relative entropy of states of a von Neumann algebra, 
                   Commun.Math.Phys.,Vol.114,pp.345-349 (1988).
%\bibitem{FK}     J.I.Fujii and E.Kamei, Relative operator entropy in noncommutative information theory, 
%                   Math.Japonica, Vol.34,pp.341-348 (1989).

\bibitem{Tho} C.J.Thompson, Inequality with applications in statistical mechanics, J.Math.Phys., Vol.6, pp.1812-1813(1965).
\bibitem{Tsa}    C.Tsallis, Possible generalization of Bolzmann-Gibbs statistics, J.Stat.Phys.,Vol.52,pp.479-487 (1988).
\bibitem{AO} C.Tsallis et al., {\em Nonextensive Statistical Mechanics and Its Applications}, edited by S. Abe and Y. Okamoto (Springer-Verlag, Heidelberg,2001); 
see also the comprehensive list of references at \texttt{http://tsallis.cat.cbpf.br/biblio.htm}.
\bibitem{WZ} B.Y.Wang and F.Zhang, Trace and eigenvalue inequalities for ordinary and Hadamard products of positive semidefinite Hermitian matrices, 
SIAM J.Matrix.Anal.Appl.,Vol.16,No.4,pp.1173-1183 (1995). 
%\bibitem{Zhan} X.Zhan, Matrix Inequalities, Springer, 2002.
%\bibitem{HP}     F.Hiai and D.Petz, The proper formula for relative entropy and its asymptotics in quantum probability, 
%                   Commun.Math.Phys.,Vol.143,pp.99-114 (1991).
%\bibitem{HP2}    F.Hiai and D.Petz, The Golden-Thompson trace inequality is complimented, Linear Alg. Appl.,Vol.181,pp.153-185 (1993).
%\bibitem{MO}     A.W.Marshall and I.Olkin, Inequalities for the trace function, Aequationes Mathematicae, Vol.29, pp.35-39 (1985).


%\bibitem{Bha}    R.Bhatia, Matrix Analysis, Springer, 1997.
%\bibitem{Zhan}   X.Zhan, Matrix inequalities, Springer, 2002.
%\bibitem{Han}    F.Hansen, An operator inequality, Math.Ann.,Vol.246, pp.249-250 (1980).





%\bibitem{Cov} T.M.Cover and J.A.Thomas, Elements of Information Theory, John Wiley and Sons, 1991.
%\bibitem{Fan} M.Fannes, A continuity property of entropy density for spin lattice systems, Commun.Math.Phys.,Vol.31,pp.291-294 (1973).
%\bibitem{AF} R.Alicki and M.Fannes, Quantum Dynamical Systems, Oxford University Press, 2001.
%\bibitem{NC} M.A.Nielsen and I.Chuang, Quantum Computation and Quantum Information, Cambridge Press,2000.
%\bibitem{An} T.Ando, Topics on Operator Inequalities, Lecture Notes, Hokkaido Univ.,Sapporo,1978.
%\bibitem{Petz} D.Petz, Quasi-entropies for finite quantum system, Rep.Math.Phys.,Vol.23,pp.57-65(1986).
\end{thebibliography}

\documentclass[11pt]{article}
\usepackage{graphicx}
\usepackage{amsfonts}
\usepackage{theorem}
\newtheorem{Lem}{Lemma}[section]
\newtheorem{Def}[Lem]{Definition}
\newtheorem{The}[Lem]{Theorem}
\newtheorem{Prop}[Lem]{Proposition}
\newtheorem{Cor}[Lem]{Corollary}

\newcommand{\qed}{\hbox{\rule{6pt}{6pt}}}

\newcommand{\Tr}{\mathbf{Tr}}

\begin{document}
\title{Trace inequalities in nonextensive statistical mechanics }
\author{Shigeru Furuichi$^1$\footnote{E-mail:furuichi@ed.yama.tus.ac.jp}\\
$^1${\small Department of Electronics and Computer Science,}\\{\small Tokyo University of Science, Yamaguchi, 756-0884, Japan}}
\date{}
\maketitle
{\bf Abstract.} In this short paper, we establish a variational expression of the Tsallis relative entropy. 
In addition, we derive a generalized thermodynamic inequality and a generalized Peierls-Bogoliubov inequality. 
Finally we give a generalized Golden-Thompson inequality.
\vspace{3mm}

{\bf Keywords : } Trace inequalities, thermodynamic inequality, Peierls-Bogoliubov inequality, Golden-Thompson inequality and Tsallis relative entropy 

\vspace{3mm}

{\bf 2000 Mathematics Subject Classification : } 47A63, 94A17, 15A39

\vspace{3mm}

%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%
%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%
%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%
%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%  Section1  %%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%
%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%
%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%
%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%
\section{Introduction}

Recently, the matrix trace inequalities in statistical mechanics are studied by Bebiano et.al. in \cite{BPL}.
 Their results are generalized by them in \cite{BPL2} via $\alpha$-power mean.
In addition, the further generalized logarithmic trace inequalities are obtained 
and their convergences are shown via generalized Lie-Trotter formulae in \cite{Furuta}.
Inspired by their works, we generalize the trace inequalities in \cite{BPL} 
by means of a parametric extended logarithmic function $\ln_{\lambda}$ which will be defined below. 
Our generalizations are different from those in \cite{BPL2,Furuta}.
In the sense of our generalization, we give a generalized Golden-Thompson inequality. In addition, we give a related trace inequality as concluding remarks.

We denote $e_{\lambda}^x \equiv (1 + \lambda x)^{\frac{1}{\lambda}}$ and its inverse function $\ln_{\lambda}x \equiv \frac{x^{\lambda} -1}{\lambda}$,
for $\lambda \in (0,1]$ and $x \geq 0$. 
The functions $e_{\lambda}^x $ and $\ln_{\lambda} x$ converge to $e^x$ and $\log x$ as $\lambda \to 0$, respectively.
Note that we have the following relations:
\begin{equation}
 e_{\lambda}^{x+y+\lambda xy} = e_{\lambda}^x e_{\lambda}^y, \,\,\,\,\,
 \ln_{\lambda} xy = \ln_{\lambda}x + \ln_{\lambda} y + \lambda \ln_{\lambda} x \ln_{\lambda} y. \label{nonex01}
\end{equation}
The Tsallis entropy was originally defined in \cite{Tsa} by $- \sum_{i=1}^n p_i^q \ln_{1-q} p_i = \frac{\sum_{i=1}^n (p_i^q - p_i)}{1-q}$ for any nonnegative real number $q$ and
a probability distribution $p_i \equiv p(X=x_i)$ of a given random variable $X$. 
Taking the limit as $q \to 1$, the Tsallis entropy converges to the Shannon entropy $-\sum_{i=1}^n p_i \log p_i$. 
We may regard that the expectation value $E_q(X) = \sum_{i=1}^n p_i^q x_i$ depending on the parameter $q$ is adopted in order to define the Tsallis entropy as $x_i =-\ln_{1-q}p_i$, 
while the usual expectation value $E(X) = \sum_{i=1}^n p_i x_i$ is adopted in order to define the Shannon entropy as $x_i =-\log p_i$.
In the sequel we use the parameter $\lambda \in (0,1]$ insead of $q$. There is a relation between these two parameters such that $q = 1-\lambda$.  

The Tsallis entropy and the Tsallis relative entropy in quantum system (noncommutative system) are defined in the following manner. 
See \cite{FYK,FYK2} for example.
\begin{Def}
The Tsallis entropy is defined by 
$$
S_{\lambda}(\rho) \equiv \frac{\hbox{Tr}[ \rho^{1-\lambda}  - \rho]}{\lambda} = -\hbox{Tr}[\rho^{1-\lambda} \ln_{\lambda} \rho]
$$
for a density operator $\rho$ and $\lambda \in (0,1]$. The Tsallis relative entropy is defined by
$$
D_{\lambda} (\rho \vert \sigma) \equiv \frac{\hbox{Tr}[\rho - \rho^{1-\lambda}\sigma^{\lambda}]}{\lambda} = \hbox{Tr}[\rho^{1-\lambda} (\ln_{\lambda} \rho - \ln_{\lambda} \sigma)]
$$
for density operators $\rho$, $\sigma$  and $\lambda \in (0,1]$.
\end{Def}
The Tsallis entropy and the Tsallis relative entropy converge to the von Neumann entropy $S(\rho) \equiv -\hbox{Tr}[\rho \log \rho]$ 
and the relative entropy $D(\rho \vert \sigma) \equiv \hbox{Tr}[\rho(\log \rho -\log \sigma)]$ as $\lambda \to 0$, respectively. 
See \cite{OP} for details on the theory of quantum entropy.  
Two Tsallis entropies have nonadditivities such that
\begin{equation}  \label{pseudo_Tsallis}
S_{\lambda}(\rho_1\otimes \rho_2) =S_{\lambda}(\rho_1)+S_{\lambda}(\rho_2)+\lambda S_{ \lambda}(\rho_1)S_{\lambda}(\rho_2),
\end{equation}
and
\begin{equation}  \label{pseudo_relative_Tsallis}
D_{ \lambda}  (\rho_1 \otimes \rho_2 \vert \sigma_1 \otimes \sigma_2) =D_{ \lambda}  (\rho_1\vert \sigma_1)+D_{ \lambda}  (\rho_2\vert \sigma_2 )
-\lambda D_{ \lambda}  (\rho_1\vert \sigma_1)D_{ \lambda}  (\rho_2\vert \sigma_2),
\end{equation}
due to the nonadditivity Eq.(\ref{nonex01}) of the function $\ln_{\lambda}$.
Thus the field of the study using these entropies is often called the nonextensive statistical physics 
and many research papers have been published in mainly statistical physics \cite{AO}. 
In addition, for the relative R\'enyi entropy of order $\lambda$ 
$$
R_{\lambda}(\rho\vert \sigma) \equiv \frac{1}{\lambda}\log Tr[\rho^{1-\lambda}\sigma^{\lambda}],\quad \lambda \in (0,1]
$$
we have the following relation between $R_{\lambda}(\rho\vert \sigma)$ and $D_{\lambda}(\rho\vert\sigma)$ :
$$
\lambda D_{\lambda}(\rho\vert\sigma)+\exp[\lambda R_{\lambda}(\rho\vert \sigma)] = 1.
$$
See our previous papers \cite{FYK,FYK2} on the mathematical properties of the Tsallis entropy and the Tsallis relative entropy.

\section{A variational expression of Tsallis relative entropy}

In this section, we derive a variational expression of the Tsallis relative entropy as a parametric extension of that of the relative entropy in Lemma 1.2 of \cite{HP2}.
A variational expression of the relative entropy has been studied in the general setting of von Neumann algebras \cite{Petz2,Kos}. 
In the sequel, we consider $n \times n$ complex matrices in the finite quantum system.
A Hermitian matrix $A$ is called a nonnegative matrix (and denoted by $A \geq 0$) if $\langle x, Ax\rangle \geq 0$ for all $x \in \hbox{C}^n$. 
A nonnegative matrix $A$ is called a positive matrix (and denoted by $A >0$)
if it is invertible. Throughout this paper, $\Tr$ means the usual matrix trace. A nonnegative matrix $A$ is called a density matrix if $\Tr[A]=1$.
In the below, we sometimes relax the condition of the unital trace for the matrices in the definition of the Tsallis relative entropy $D_{\lambda}(\cdot \vert \cdot)$, 
since it is not essential in the mathematical studies of the entropic functionals.
\begin{The}   \label{gen_rela_vari_exp}
For $\lambda \in (0,1]$, we have the following relations.
\begin{itemize}
\item[(1)] If $A$ and $Y$ are nonnegative matrices, then
$$ \ln_{\lambda} \Tr[ e_{\lambda}^{A+\ln_{\lambda}Y }] = \max \left\{ \Tr[X^{1-\lambda}A] -D_{\lambda} (X\vert Y) : X \geq 0, \Tr[X] =1 \right\}. $$
\item[(2)] If $X$ is a positive matrix with $\Tr[X] = 1$ and $B$ is a Hermitian matrix, then
$$  D_{\lambda}(X\vert e_{\lambda}^B)  =\max \left\{ \Tr[X^{1-\lambda} A] - \ln_{\lambda} \Tr[e_{\lambda}^{A+B}] : A \geq 0 \right\}.$$
\end{itemize}
\end{The}

{\it Proof}:
The proof is almost similar to that of Lemma 1.2 in \cite{HP2}.
\begin{itemize}
\item[(1)] 
For $\lambda = 1$, we have $\ln_{\lambda} \Tr[ e_{\lambda}^{A+\ln_{\lambda}Y }] =  \Tr[X^{1-\lambda}A] -D_{\lambda} (X\vert Y) $ by easy calculations with the usual convention $X^0=I$. 
Thus we assume $\lambda \in (0,1)$.
Let us denote
$$F_{\lambda}(X) = \Tr[X^{1-\lambda} A] -D_{\lambda} (X\vert Y) $$
for a nonnegative matrix $X$ with $\Tr[X]=1$. If we take the Schatten decomposition $X = \sum_{j=1}^n r_j E_j $, where all
$E_j$, $(j=1,2,\cdots ,n)$ are projections of rank one with $\sum_{j=1}^n E_j =I$ and $r_j \geq 0$, $(j=1,2,\cdots ,n)$ with $\sum_{j=1}^n r_j =1$, then we rewrite
$$F_{\lambda}\left(\sum_{j=1}^n r_j E_j\right) = \sum_{j=1}^n \left\{ r_j^{1-\lambda}\Tr[E_j A] + \frac{1}{\lambda} r_j^{1-\lambda}  \Tr[E_jY^{\lambda}] - \frac{1}{\lambda} r_j \Tr[E_j]\right\}.$$
Since we have 
$$ \frac{\partial}{\partial r_j} F_{\lambda} \left( \sum_{j=1}^n r_j E_j\right) \vert _{r_j =0} = +\infty$$ 
%(which means $F_{\lambda}$ is concave function),  because $f$: concave iff $f(x) \leq f(c) +f'(c)(x-c),\,\, (a\leq x \leq b)$),
$F_{\lambda}(X)$ attains its maximum at a nonnegative matrix $X_0$ with $\Tr[X_0] = 1$. Then for any Hermitian matrix $S$ with $\Tr[S] =0$ (since $\Tr[X_0 + t S]$ must be $1$), we have
$$0 = \frac{d}{dt} F_{\lambda}(X_0+tS) \vert _{t=0} = (1-\lambda) \Tr[S (X_0^{-\lambda}A +\frac{1}{\lambda} X_0^{-\lambda}Y^{\lambda} )],$$
 so that $X_0^{-\lambda}A +\frac{1}{\lambda} X_0^{-\lambda}Y^{\lambda} = c I$ for $c \in \mathbf{R}$. Thus we have 
$$ X_0 = \frac{e_{\lambda}^{A+\ln_{\lambda}Y}}{\Tr[e_{\lambda}^{A+\ln_{\lambda}Y}]}$$  
by the normalization condition. By the formulae $\ln_{\lambda}\frac{y}{x} = \ln_{\lambda}y + y^{\lambda} \ln_{\lambda} \frac{1}{x}$ 
and $\ln_{\lambda} \frac{1}{x} = -x^{-\lambda} \ln_{\lambda}x$, we have
$$F_{\lambda}(X_0) = \ln_{\lambda}\Tr[e_{\lambda}^{A+\ln_{\lambda}Y}]. $$
\item[(2)] For $\lambda = 1$, we have $D_{\lambda}(X\vert e_{\lambda}^B)  =  \Tr[X^{1-\lambda} A] - \ln_{\lambda} \Tr[e_{\lambda}^{A+B}] $ 
 by easy calculations. Thus we assume $\lambda \in (0,1)$.
It follows from (1) that the functional $g(A) = \ln_{\lambda} \Tr[e_{\lambda}^{A+B}]$ 
defined on the set of all nonnegative matrices is convex, due to triangle inequality on $\max$. 
Now let $A_0 = \ln_{\lambda} X -B$, and denote
$$G_{\lambda}(A) = \Tr[X^{1-\lambda}A] - \ln_{\lambda}\Tr[e_{\lambda}^{A+B}], $$
which is concave on the set of all nonnegative matrices. Then for any nonnegative matrix $S$, there exists a nonnegative matrix $A_0$ such that
$$\frac{d}{dt}G_{\lambda}(A_0+tS) \vert _{t=0} = 0. $$
Therefore $G_{\lambda}(A)$ attaines the maximum $G_{\lambda}(A_0) = D_{\lambda}(X\vert e_{\lambda}^B)$.
\end{itemize}
\hfill \qed
\vspace*{2mm}

Taking the limit as $\lambda \to 0$, Theorem \ref{gen_rela_vari_exp} recovers the similar form of Lemma 1.2 in \cite{HP2} under the assumption of nonnegativitiy of $A$.
If $Y=I$ and $B=0$ in (1) and (2) of Theorem \ref{gen_rela_vari_exp}, respectively, then we obtain the following corollary.

\begin{Cor}\label{gen_vari_exp}
\begin{itemize}
\item[(1)] If $A$ is a nonnegative matrix, then
$$\ln_{\lambda}\Tr[e_{\lambda}^A] = \max \left\{ \Tr[X^{1-\lambda}A] + S_{\lambda}(X) : X \geq 0,\, \Tr[X]=1\right\}.$$
\item[(2)] For a density matrix $X$, we have
$$-S_{\lambda}(X) = \max \left\{ \Tr[X^{1-\lambda}A] -\ln_{\lambda}\Tr[e_{\lambda}^A] : A \geq 0 \right\}. $$
\end{itemize}
\end{Cor}
Taking the limit as $\lambda \to 0$, Corollary \ref{gen_vari_exp} recovers the similar form of Theorem 1 in \cite{BPL} under the assumption of nonnegativitiy of $A$.

%\begin{Rem}
%Form (1) of Corollary \ref{gen_vari_exp}, we may regard the functional $f_{\lambda} (X) \equiv \Tr[X^{1-\lambda}A] + S_{\lambda}(X)$
%depending on $\lambda \in (0,1]$ 
%as a generalized free energy functional of the quantum state $X$.
%\end{Rem}

\section{Generalized logarithmic trace inequalities}
In this section, we derive some trace inequalities in terms of the results obtained in the previous section.
From (1) of Corollary \ref{gen_vari_exp}, we have the generalized thermodynamic inequality:
\begin{equation}  \label{gen_Thermo_ineq}
\ln_{\lambda}\Tr[e_{\lambda}^H] \geq \Tr[D^{1-\lambda}H] + S_{\lambda}(D), 
\end{equation}
for a density matrix $D$ and a nonnegative matrix $H$.
Putting $D = \frac{A}{\Tr[A]}$ and $H = \ln_{\lambda}B$ in Eq.(\ref{gen_Thermo_ineq}) for $A \geq 0$ and $B \geq I$, 
we have the generalized Peierls-Bogoliubov inequality (cf.Theorem 3.3 of \cite{FYK}):
\begin{equation}  \label{gen_Pei_Bog}
(\Tr[A])^{1-\lambda} \left( \ln_{\lambda}\Tr[A]  - \ln_{\lambda} \Tr[B] \right) \leq  \Tr[A^{1-\lambda} (\ln_{\lambda}A - \ln_{\lambda} B) ],
\end{equation}
for nonnegative matrices $A$ and $B \geq I$.

\begin{Lem} \label{vari_exp_rela_lemma}
The following statements are equivalent.
\begin{itemize}
\item[(1)] $F_{\lambda}(A) = \ln_{\lambda} \Tr[e_{\lambda}^A]$ is convex in a Hermitian matrix $A$.
\item[(2)] $f_{\lambda}(t) = \ln_{\lambda} \Tr[e_{\lambda}^{A+tB}] $ is convex in $t\in \hbox{R}$.
\end{itemize}
\end{Lem}
{\it Proof}:
 Putting $A_1 = A+ x B$ and $A_2 = A + y B$ in
$$\ln_{\lambda} \Tr[e_{\lambda}^{\mu A_1+(1-\mu_2) A_2}] \leq \mu \ln_{\lambda} \Tr[e_{\lambda}^{A_1}] +(1-\mu) \ln_{\lambda}\Tr[e_{\lambda}^{A_2}], $$
we find the convexity of $f_{\lambda}(t)$. Thus (1) implies (2).

Conversely, putting $A = A_2$, $B= A_1 - A_2$ and $x=1,y=0$ in
$$\ln_{\lambda}\Tr[e_{\lambda}^{A + (\mu x +(1-\mu)y )B}] \leq \mu \ln_{\lambda}\Tr[e_{\lambda}^{A + x B}] + (1-\mu) \ln_{\lambda} \Tr[e_{\lambda}^{A + yB}], $$
we find the convexity of $F_{\lambda}(A)$. Thus (2) implies (1).

\hfill \qed
\vspace*{2mm}

\begin{Cor}  \label{corollary_trace_inequality}
For nonnegative matrices $A$ and $B$, we have
\begin{equation}  \label{cor_from_bogo_ineq}
\ln_{\lambda}\Tr[e_{\lambda}^{A+B}] - \ln_{\lambda}\Tr[e_{\lambda}^A] \geq \frac{\Tr[B (e_{\lambda}^A)^{1-\lambda}]}{(\Tr[e_{\lambda}^A])^{1-\lambda}}. 
\end{equation}
\end{Cor}
{\it Proof}:
From Theorem \ref{gen_rela_vari_exp}, $F_{\lambda}(A)=\ln_{\lambda}\Tr[e_{\lambda}^A]$ is convex. Due to Lemma \ref{vari_exp_rela_lemma},
$f_{\lambda}(t) = \ln_{\lambda} \Tr[e_{\lambda}^{A+tB}]$ is convex, which implies that $f_{\lambda}(1) \geq f_{\lambda}(0) +f'_{\lambda}(0)$. 
Thus we have the corollary by simple calculations.  
\hfill \qed
\vspace*{2mm}

%Note that we also obtain the inequality (\ref{cor_from_bogo_ineq}) by putting $A = e_{\lambda}^H$ and $B = e_{\lambda}^{H+K}$ in the inequality (\ref{gen_Pei_Bog}).
%In addition, putting $A=\ln_{\lambda}D$ and $B= H -\ln_{\lambda}D$ for a density matrix $D$ in the inequality (\ref{cor_from_bogo_ineq}),
% we obtain the inequality (\ref{gen_Thermo_ineq}).
%Thus we have the following proposition.

%\begin{Prop}  \label{prop_tra_ineq_gen}
%The following conditions are equivalent.
%\begin{itemize}
%\item[(1)] The generalized thermodynamic inequality (\ref{gen_Thermo_ineq}).
%\item[(2)] The generalized Peierls-Bogoliubov inequality (\ref{gen_Pei_Bog}).
%\item[(3)] The trace inequality (\ref{cor_from_bogo_ineq}) given in Corollary \ref{corollary_trace_inequality}.
%\end{itemize}
%\end{Prop}
%Taking the limit as $\lambda \to 0$, Proposition \ref{prop_tra_ineq_gen} recovers Theorem 2 in \cite{BPL}.

\section{Generalized exponential trace inequality}
For nonnegative real numbers $x,y$ and $0< \lambda \leq 1$, the relations $e_{\lambda}^{x+y} \leq e_{\lambda}^{x+y+\lambda xy} =e_{\lambda}^{x} e_{\lambda}^{y}$ hold.
These relations naturally motivate us to consider the following inequalities in the noncommutative case.

\begin{Prop} \label{GT_ineq_gen_before}
For nonnegative matrices $X$ and $Y$, and $0 < \lambda \leq 1$, we have
$$
\Tr[  e_{\lambda}^{X+Y}  ] \leq \Tr[  e_{\lambda}^{X+Y+\lambda Y^{1/2}XY^{1/2}}  ].
$$
\end{Prop}
{\bf (Proof)}
 Since $Y^{1/2}XY^{1/2} \geq 0$ for nonnegative matrices $X$ and $Y$, we have
$$ 
I+\lambda (X+Y) \leq I +\lambda \left\{X+Y + \lambda (Y^{1/2}XY^{1/2})\right\}.
$$
It is known \cite{Petz1} that $A \leq B$ implies $\Tr[f(A)] \leq \Tr[f(B)]$, if $A$ and $B$ are Hermitian matrices and $f:\hbox{R} \to \hbox{R}$ is an increasing function.
Thus we have 
$$
\Tr[\left(I+\lambda (X+Y)\right)^{1/\lambda}] \leq \Tr[\left(I +\lambda \left\{X+Y + \lambda (Y^{1/2}XY^{1/2})\right\}\right)^{1/\lambda}].
$$
for $0 < \lambda \leq 1$.
\hfill \qed 
\vspace*{2mm}

Note that we have the matrix inequality :
$$
e_{\lambda}^{X+Y}  \leq e_{\lambda}^{X+Y+\lambda Y^{1/2}XY^{1/2}}  
$$
for $\lambda \geq 1$ by the application of the L\"owner-Heinz inequality \cite{Low,Hei,Ped}.

\begin{Prop}  \label{GT_ineq_gen}
For nonnegative matrices $X,Y$, and $\lambda \in (0,1]$, we have
\begin{equation}   \label{GT_ineq_gen_EQ}
\Tr[e_{\lambda}^{X+Y+\lambda XY}] \leq \Tr[e_{\lambda}^X e_{\lambda}^Y].
\end{equation}
\end{Prop}
{\it Proof}:
For $\lambda \in (0,1] $, we have 
\begin{equation}  \label{mod_lieb_thirring}
\Tr[(AB)^{1/\lambda}] \leq \Tr[A^{1/\lambda} B^{1/\lambda}]
\end{equation}
from the special case of the Lieb-Thirring inequality \cite{LT} (see also \cite{HH,WZ}), 
$\Tr[(AB)^{\alpha}] \leq \Tr[A^{\alpha}B^{\alpha}]$ for nonnegative matrices $A$, $B$ and any real number $\alpha \geq 1$.
Putting $A= I +\lambda X$ and $B=I +\lambda Y$ in Eq.(\ref{mod_lieb_thirring}), 
we have 
$$
\Tr[\left\{ (I +\lambda X)(I +\lambda Y)\right\}^{1/\lambda}]  \leq \Tr[  (I +\lambda X)^{1/\lambda} (I +\lambda Y)^{1/\lambda}],
$$
which is the desired one.
\hfill \qed 
\vspace*{2mm}

Notice that Golden-Thompson inequality \cite{Gol,Tho}, 
$$\Tr[e^{X+Y}] \leq \Tr[e^X e^Y]$$
which holds for Hermitian matrices $X$ and $Y$, is recovered by taking the limit as $\lambda \to 0$ in Proposition \ref{GT_ineq_gen},
in particular case of nonnegative matrices $X$ and $Y$.

\section{Concluding remarks}

Since $\Tr[HZHZ] \leq \Tr[H^2Z^2]$ for Hermitian matrices $H$ and $Z$ \cite{Petz2,Fujii}, we have for nonnegative matrices $X$ and $Y$,
$$\Tr[(I+X+Y+Y^{1/2}XY^{1/2})^2] \leq \Tr[(I+X+Y+XY)^2]$$
 by easy calculations. This implies the inequality
$$ \Tr[e_{1/2}^{X+Y+1/2Y^{1/2}XY^{1/2}}] \leq \Tr[e_{1/2}^{X+Y+1/2XY}].$$
 Thus we have 
\begin{equation}   \label{eq_1_2_GT_gen}
\Tr[e_{1/2}^{X+Y}] \leq \Tr[e_{1/2}^Xe_{1/2}^Y]
\end{equation}
from Proposition \ref{GT_ineq_gen_before} and Proposition \ref{GT_ineq_gen}. 
Putting $B=\ln_{1/2}Y$ and $A=\ln_{1/2}Y^{-1/2}XY^{-1/2}$ in (2) of Theorem \ref{gen_rela_vari_exp} 
under the assumption of $I\leq Y\leq X$ and using Eq.(\ref{eq_1_2_GT_gen}), 
we have
\begin{eqnarray}
D_{1/2}(X\vert Y)  &=& D_{1/2}(X\vert e_{1/2}^{\ln_{1/2}Y}) \nonumber \\
&\geq& \Tr[X^{1/2}A] -\ln_{1/2} \Tr[e_{1/2}^{A+B}] \nonumber \\
&\geq&  \Tr[X^{1/2}A] -\ln_{1/2} \Tr[  e_{1/2}^{A} e_{1/2}^{B}    ] \nonumber \\
&=& \Tr[X^{1/2}\ln_{1/2} Y^{-1/2}XY^{-1/2} ]  -\ln_{1/2} \Tr[Y^{-1/2}XY^{-1/2} Y] \nonumber \\
&=& \Tr[X^{1/2}\ln_{1/2} Y^{-1/2}XY^{-1/2} ],  \label{lower_bound_rel_ent_1_2}
\end{eqnarray}
which gives a lower bound of the Tsallis relative entropy in the case of $\lambda = 1/2$ and $I\leq Y\leq X$.

\section*{Acknowledgement}
We would like to thank the reviewers for providing valuable comments to improve the manuscript.
This work was supported by the Japanese Ministry of Education, Science, Sports and Culture, Grant-in-Aid for 
Encouragement of Young scientists (B), 17740068.

\end{document}